OpenStudy (anonymous):

For 2x – 3y = –10, when x = 4, What is the value of y?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Plug in x as 4 2(4) - 3y = -10 8 - 3y = -10 -3y = -10-8 -3y = -18 y = 6

OpenStudy (anonymous):

oh ok thanks